OpenStudy (anonymous):

what is the missing value in the ? 16x^2+?x+36

OpenStudy (anonymous):

what is 16 the square of?

OpenStudy (anonymous):

would it be 8 or 6

OpenStudy (anonymous):

neither what number squared it 16?

OpenStudy (anonymous):

4

OpenStudy (anonymous):

4x

OpenStudy (anonymous):

ok good what number squared is 36?

OpenStudy (anonymous):

6

OpenStudy (anonymous):

fine what is \(2\times 4\times 6\) ?

OpenStudy (anonymous):

56

OpenStudy (anonymous):

hmmm

OpenStudy (anonymous):

i meant 48 xD my bad

OpenStudy (anonymous):

lol ok that is your answer

OpenStudy (anonymous):

\[16x^2+48x+36=(4x+6)^2\]

OpenStudy (anonymous):

these problems always have the same pattern, that you are being taught (more or less) to recognize \[(a+b)^2=a^2+2ab+b^2\]